OpenStudy (goldrush18):

a.find the gradient of the curve f(x)=x^2-4x+3 at x=3 b. find the first derivative of the function y=2-3x^2 at the point (1, -1)

OpenStudy (anonymous):

For Slope find f'(x)

OpenStudy (anonymous):

f(x)=x^2-4x+3 f'(x) = 2x - 4

OpenStudy (anonymous):

Nw put x = 3 f'(x) = 2* 3 - 4 = 2

OpenStudy (abb0t):

Follow @Yahoo! he is givin away words of wisdom :)

OpenStudy (goldrush18):

thank u so much :)

OpenStudy (goldrush18):

but what about part b @Yahoo! what must i do with the 2 points

OpenStudy (anonymous):

Just Differentiate